# UniHGKR: Unified Instruction-aware Heterogeneous Knowledge Retrievers

### v1

@@ Abstract
Existing information retrieval (IR) models often assume a homogeneous structure for knowledge sources and user queries, limiting their applicability in real-world settings where retrieval is inherently heterogeneous and diverse. In this paper, we introduce UniHGKR, a unified instruction-aware heterogeneous knowledge retriever that (1) builds a unified retrieval space for heterogeneous knowledge and (2) follows diverse user instructions to retrieve knowledge of specified types. UniHGKR consists of three principal stages: heterogeneous self-supervised pretraining, text-anchored embedding alignment, and instruction-aware retriever fine-tuning, enabling it to generalize across varied retrieval contexts. This framework is highly scalable, with a BERT-based version and a UniHGKR-7B version trained on large language models. Also, we introduce CompMix-IR, the first native heterogeneous knowledge retrieval benchmark. It includes two retrieval scenarios with various instructions, over 9,400 question-answer (QA) pairs, and a corpus of 10 million entries, covering four different types of data. Extensive experiments show that UniHGKR consistently outperforms state-of-the-art methods on CompMix-IR, achieving up to 6.36% and 54.23% relative improvements in two scenarios, respectively. Finally, by equipping our retriever for open-domain heterogeneous QA systems, we achieve a new state-of-the-art result on the popular ConvMix Christmann et al. (2022b) task, with an absolute improvement of up to 4.80 points. [^1] Our code, datasets and model checkpoints are available at: https://github.com/ZhishanQ/UniHGKR
@@
Dehai Min1, Zhiyang Xu3, Guilin Qi1, Lifu Huang4, Chenyu You2 1Southeast University, 2Stony Brook University, 3Virginia Tech, 4UC Davis
@@ 1 Introduction
Retrieval-Augmented Generation (RAG Gao et al. (2023)) has become a pivotal technique for improving the faithfulness of generative large language models (LLMs Achiam et al. (2023)). By leveraging retrievers to extract relevant knowledge from large-scale knowledge corpus, RAG effectively reduces the hallucinations often produced by LLMs Ayala and Bechard (2024).
@@
In addition, existing heterogeneous IR benchmarks have limited knowledge coverage Petroni et al. (2021); Muennighoff et al. (2023). For example, studies like Chen et al. (2021); Zhong et al. (2022) focus only on two types of knowledge: tables and text. To address this gap, we introduce  CompMix-IR, the first-ever benchmark for heterogeneous knowledge retrieval. CompMix-IR has over 9,400 QA pairs and a corpus of 10 million entries spanning four distinct knowledge types: Text, Knowledge Graphs (KG), Tables, and Infoboxes. Derived from the open-domain QA dataset CompMix Christmann et al. (2024), CompMix-IR transforms this QA task into a standard IR task (as detailed in Section 3). To better reflect real-world retrieval needs, we define two distinct scenarios in this benchmark: (1) retrieving relevant evidence across all knowledge types, and (2) retrieving evidence of a specific type, as specified by user instructions. Both scenarios utilize the same evidence pool, requiring the retriever to adapt query-evidence similarity based on the instructions. This setup mirrors the complexities of real-world retrieval tasks, offering enhanced practical relevance and utility for diverse applications.
Experimental results demonstrate the effectiveness of our proposed UniHGKR over the existing methods, with relative improvements of up to 6.36% and 54.23% in two different scenarios. In addition to the BERT-based UniHGKR-base model, we also extend our framework to an LLM-based retriever and train the UniHGKR-7B model to verify scalability. Both models achieve state-of-the-art (SOTA) performance on CompMix-IR respective to their parameter scales. Furthermore, in the context of open-domain heterogeneous QA, systems equipped with UniHGKR retriever set a new SOTA on the ConvMix task Christmann et al. (2022b), with an absolute gain of up to 4.80 points, further validating its real-world applicability.
@@ 2 Related Work
QA over Heterogeneous Knowledge. Each data type has its own characteristics and provides unique benefits. Some studies explores the integration of knowledge sources to QA Ma et al. (2022a); Min et al. (2024). For instance, HybridQA Chen et al. (2020b) and OTT-QA Chen et al. (2021) investigate the task of extracting answers from the combination of tables and text. Going further, CONVINSE Christmann et al. (2022b), Explaignn Christmann et al. (2023) and FAITH Jia et al. (2024) consider four knowledge sources like this paper. However, their primary focus is on the answer generation parts of the system. Their retrieval approach is a time-consuming online pipeline: identifying entity IDs in questions, then conducting online searches in Wikipedia and Wikidata Vrandečić and Krötzsch (2014), and finally employing BM25 Robertson et al. (2009) to rank a small set of evidence.
3 CompMix-IR Benchmark
In this section, we provide a detailed description about the CompMix-IR and the instruction schema.
3.1 Heterogeneous Knowledge Collection
  • KG facts. We use CLOCQ Christmann et al. (2022a) to retrieve the top-1000 KG triples related to each question from the Wikidata dump. The retrieved KG facts are then linearized, with entities and relations separated by commas.
  • Text, Tables and Infoboxes. We use the IDs of entities mentioned in questions to retrieve the corresponding Wikipedia pages. Subsequently, a parser is used to extracts natural language paragraphs (text evidence), tables, and infoboxes from the pages. Following Oguz et al. (2022), both tables and infoboxes are linearized using simple templates, and page titles added at the head of evidence for clearer information.
Types Avg. length Count Percentage
Text 19.86 5,916,596 57.74%
KG 11.40 2,214,854 21.61%
Table 20.32 1,043,105 10.18%
Infobox 11.05 1,072,440 10.47%
Sum 17.18 10,246,995 100.00%
To align with the standard IR task setup, we use automated scripts to label relevant evidence (golden labels) for each question. The labeling criteria require that the evidence contains entities mentioned in both the question and answer text. Each question has at least one piece of evidence that provides the answer. The evidence retrieved for all questions in CompMix is combined into a heterogeneous knowledge pool, forming the corpus for the IR task. This corpus includes over 10 million pieces of evidence, covering knowledge about 137,808 different entities. Detailed statistics of CompMix-IR are presented in Table 2, and examples of evidence and QA pair, as well as additional statistics, provided in Appendix A.
@@ 5.1 Baselines
Fine-tuned Baselines. We follow the verbalizer-retriever approach from UDT-QA Ma et al. (2022b) to fine-tune a BERT-base model, serving as the UDT retriever. Since UDT focuses on homogeneous textual representations of heterogeneous data, we replace $d_{i}$ with $t_{i}$ from the data-text pairs $\mathcal{D}$ during its training and evaluation, ensuring this model only interacts with the textual corpus. For comparison, we also fine-tune a BERT-base model on the original CompMix-IR. Additionally, we finetune a DPR model using the UniK-QA method Oguz et al. (2022), serving as the UniK retriever. All fine-tuning uses the same positive and hard negative samples as UniHGKR. For baseline models lacking instruction-following capabilities, we input only the query across all retrieval scenarios to ensure optimal performance.
@@ 5.2 Evaluation Metrics
Retrieval Scenario 1 (instruction $\bm{I}_{\text{All}}$) Retrieval Scenario 2 (instruction $\bm{I}_{\tau}$) Method Size Ins Hit@5 Hit@10 Hit@100 MRR@100 KG-Hit Text-Hit Table-Hit Info-Hit BM25 - 11.51 17.40 52.39 8.54 24.2 34.55 8.5 19.79 DPR 109M 24.89 36.32 78.76 17.51 49.13 63.68 15.63 41.57 Mpnet 109M 26.23 37.99 82.67 18.46 63.02 61.11 18.96 52.1 GTR-T5-base 110M 24.46 36.54 80.32 16.73 57.78 59.8 22.87 46.09 Contriever 109M 28.58 40.70 83.79 20.07 62.26 63.86 18.63 55.64 SimLM 109M 25.11 37.08 80.61 17.68 59.59 59.01 17.69 52.06 Instructor-base 110M 24.86 36.22 81.55 17.80 65.63 50.25 16.82 53.36 Instructor-large 336M 25.98 36.87 81.51 18.54 68.78 44.61 17.11 53.98 BGE 109M 26.66 39.04 84.15 19.40 68.42 57.96 22.58 56.58 BERT-finetuned 109M 24.46 35.38 78.51 17.04 57.63 54.67 17.55 48.41 UDT-retriever 109M 24.96 35.49 76.52 18.24 66.10 62.48 25.90 57.05 UniK-retriever 109M 30.68 43.42 85.20 21.22 67.40 63.21 26.74 56.04 UniHGKR-base 109M 32.38 45.55 85.75 22.57 75.43 70.30 41.24 66.21 Relative gain +5.54% +4.91% +0.65% +6.36% +9.67% +10.08% +54.23% +16.06%
Table 3: the experimental results for the two retrieval scenarios on CompMix-IR. The relative gain is calculated based on the performance of UniHGKR-base compared to the best baseline, highlighted in light orange.
@@ 6.1 Main Results
Table 3 presents the retrieval performance of various models on the CompMix-IR test set. Our UniHGKR model outperforms all baselines in both scenarios, with a maximum relative improvement of 6.36% in scenario 1 and 54.23% in scenario 2, demonstrating its effectiveness in heterogeneous knowledge retrieval. Notably, powerful open-source retrievers like BGE (trained on over 200 million high-quality text pairs) only achieve an MRR@100 below 20.0 in scenario 1 and a Table-Hit of 22.58 in scenario 2, highlighting the challenges of our constructed benchmark. While UDT-retriever, which converts heterogeneous knowledge to natural language text, shows significant gains in scenario 2 (thanks to GPT-4o-mini), it shows little improvement in scenario 1 compared with its counterpart model: BERT-finetuned. Moreover, UniK-retriever, fine-tuned the DPR model on CompMix-IR, performs well across several metrics but is suboptimal on structured data (like Table and Infobox) in scenario 2. In contrast, our UniHGKR shows the greatest improvements on metrics where existing methods struggle, particularly in retrieving structured knowledge in scenario 2. This indicates that our three-stage training approach not only creates an effective representation space for heterogeneous knowledge retrieval but also excels at following diverse user instructions.
@@ 6.2 Ablation Study
Retrieval Scenario 1 (instruction $\bm{I}_{\text{All}}$) Retrieval Scenario 2 (instruction $\bm{I}_{\tau}$) Method Hit@5 Hit@10 Hit@100 MRR@100 KG-Hit Text-Hit Table-Hit Info-Hit UniHGKR-base 32.38 45.55 85.75 22.57 75.43 70.30 41.24 66.21 w/o training stage 1 (pretrain) 29.78$\downarrow$ 42.80 $\downarrow$ 84.88 $\downarrow$ 21.54 $\downarrow$ 72.97 68.6 34.01 60.06 w/o training stage 2 (alignment) 31.41 45.02 85.14 21.92 74.75 70.01 37.99 66.04 In the training stage 3 (finetune)   w/o type-preferred loss $\mathcal{L}_{\text{preferred}}$ 31.77 44.97 85.24 22.27 73.26 69.90 33.86 61.22   w/o instructions and $\mathcal{L}_{\text{preferred}}$ 31.98 44.39 85.24 22.18 68.20 $\downarrow$ 65.59$\downarrow$ 29.20 $\downarrow$ 58.65 $\downarrow$
@@
In this subsection, we conduct ablation studies to examine the roles of different training stages and components in UniHGKR for heterogeneous knowledge retrieval. Table 4 presents the performance of various UniHGKR variants, obtained by removing either Training Stage 1 or 2, as well as excluding specific components from Stage 3. Results show that removing any stage or component leads to a significant drop in performance. For retrieval scenario 1, Stage 1 (Unified Embedding Self-Supervised Pretraining) is crucial, while for scenario 2, both designed instructions and instruction-aware type-preferred loss $\mathcal{L}_{\text{preferred}}$ are key. Removing them will result in a performance drop of up to 12.14 points in the Table-Hit metric.
Retrieval Scenario 1 (instruction $\bm{I}_{\text{All}}$) Retrieval Scenario 2 (instruction $\bm{I}_{\tau}$) Method Hit@5 Hit@10 Hit@100 MRR@100 KG-Hit Text-Hit Table-Hit Info-Hit UniHGKR-base 32.38 45.55 85.75 22.57 75.43 70.30 41.24 66.21 E5-mistral-7B 31.3 43.49 83.36 22.97 69.03 41.46 33.03 62.92 LLARA-passage 37.45 51.59 86.61 26.11 68.23 70.48 37.88 60.64 LLARA-finetuned 42.19 55.35 87.81 30.83 74.38 69.86 36.40 64.40 UniHGKR-7B 49.82 59.23 88.21 38.20 76.59 71.60 37.52 66.71 Relative gain +18.08% +7.01% +0.46% +23.91% +2.97% +1.59% -0.95% +3.59%
Table 5: Retrieval performances of UniHGKR-7B and LLM-based retrievers baselines.

### v2

@@ Abstract
Existing information retrieval (IR) models often assume a homogeneous structure for knowledge sources and user queries, limiting their applicability in real-world settings where retrieval is inherently heterogeneous and diverse. In this paper, we introduce UniHGKR, a unified instruction-aware heterogeneous knowledge retriever that (1) builds a unified retrieval space for heterogeneous knowledge and (2) follows diverse user instructions to retrieve knowledge of specified types. UniHGKR consists of three principal stages: heterogeneous self-supervised pretraining, text-anchored embedding alignment, and instruction-aware retriever fine-tuning, enabling it to generalize across varied retrieval contexts. This framework is highly scalable, with a BERT-based version and a UniHGKR-7B version trained on large language models. Also, we introduce CompMix-IR, the first native heterogeneous knowledge retrieval benchmark. It includes two retrieval scenarios with various instructions, over 9,400 question-answer (QA) pairs, and a corpus of 10 million entries, covering four different types of data. Extensive experiments show that UniHGKR consistently outperforms state-of-the-art methods on CompMix-IR, achieving up to 6.36% and 54.23% relative improvements in two scenarios, respectively. Finally, by equipping our retriever for open-domain heterogeneous QA systems, we achieve a new state-of-the-art result on the popular ConvMix Christmann et al. (2022b) task, with an absolute improvement of up to 5.90 points. [^1] Our code, datasets and model checkpoints are available at: https://github.com/ZhishanQ/UniHGKR
@@
Dehai Min1, Zhiyang Xu3, Guilin Qi1, Lifu Huang4, Chenyu You2 1Southeast University, 2Stony Brook University, 3Virginia Tech, 4UC Davis qieqiemin@gmail.com, zhiyangx@vt.edu, gqi@seu.edu.cn lfuhuang@ucdavis.edu, chenyu.you@stonybrook.edu
@@ 1 Introduction
Retrieval-Augmented Generation (RAG Lewis et al. (2020); Gao et al. (2023); Qi et al. (2024)) has become a pivotal technique for improving the faithfulness of generative large language models (LLMs Achiam et al. (2023)). By leveraging retrievers to extract relevant knowledge from large-scale knowledge corpus, RAG effectively reduces the hallucinations often produced by LLMs Ayala and Bechard (2024); Muennighoff et al. (2024).
@@
In addition, existing heterogeneous IR benchmarks have limited knowledge coverage Petroni et al. (2021); Muennighoff et al. (2023). For example, studies like Chen et al. (2021b); Zhong et al. (2022) focus only on two types of knowledge: tables and text. To address this gap, we introduce  CompMix-IR, the first-ever benchmark for heterogeneous knowledge retrieval. CompMix-IR has over 9,400 QA pairs and a corpus of 10 million entries spanning four distinct knowledge types: Text, Knowledge Graphs (KG), Tables, and Infoboxes. Derived from the open-domain QA dataset CompMix Christmann et al. (2024), CompMix-IR transforms this QA task into a standard IR task (as detailed in Section 3). To better reflect real-world retrieval needs, we define two distinct scenarios in this benchmark: (1) retrieving relevant evidence across all knowledge types, and (2) retrieving evidence of a specific type, as specified by user instructions. Both scenarios utilize the same evidence pool, requiring the retriever to adapt query-evidence similarity based on the instructions. This setup mirrors the complexities of real-world retrieval tasks, offering enhanced practical relevance and utility for diverse applications.
Experimental results demonstrate the effectiveness of our proposed UniHGKR over the existing methods, with relative improvements of up to 6.36% and 54.23% in two different scenarios. In addition to the BERT-based UniHGKR-base model, we also extend our framework to an LLM-based retriever and train the UniHGKR-7B model to verify scalability. Both models achieve state-of-the-art (SOTA) performance on CompMix-IR respective to their parameter scales. Furthermore, in the context of open-domain heterogeneous QA, systems equipped with UniHGKR retriever set a new SOTA on the ConvMix task Christmann et al. (2022b), with an absolute gain of up to 5.90 points, further validating its real-world applicability.
@@ 2 Related Work
QA over Heterogeneous Knowledge. Each data type has its own characteristics and provides unique benefits. Some studies explores the integration of knowledge sources to QA Ma et al. (2022a); Min et al. (2024); You et al. (2020a, b, 2021c, 2021a, 2021b); Chen et al. (2021a); You et al. (2022). For instance, HybridQA Chen et al. (2020b) and OTT-QA Chen et al. (2021b) investigate the task of extracting answers from the combination of tables and text. Going further, CONVINSE Christmann et al. (2022b), Explaignn Christmann et al. (2023) and FAITH Jia et al. (2024) consider four knowledge sources like this paper. However, their primary focus is on the answer generation parts of the system. Their retrieval approach is a time-consuming online pipeline: identifying entity IDs in questions, then conducting online searches in Wikipedia and Wikidata Vrandečić and Krötzsch (2014), and finally employing BM25 Robertson et al. (2009) to rank a small set of evidence.
3 CompMix-IR Benchmark
In this section, we provide a detailed description of the construction of CompMix-IR, the definition of retrieval scenarios, and their instruction schema.
3.1 Heterogeneous Knowledge Collection
  • KG facts. We use CLOCQ Christmann et al. (2022a) to retrieve the top-1000 KG triples related to each question from the Wikidata dump. We also store the disambiguations and wikidata entities information returned by CLOCQ. This information helps us evaluate the relevance between the evidence and the question. To feed the structured data into the language model, the retrieved KG facts are linearized, with entities and relations separated by commas.
  • Text, Tables and Infoboxes. We use the entities mentioned in questions to retrieve the corresponding Wikipedia pages. Subsequently, a parser is used to extracts natural language paragraphs (text evidence), tables, and infoboxes from the pages. Also, we utilize hyperlinks from Wikipedia pages to map the corresponding entity mentions to Wikidata IDs. This achieves the same labeling format as KG evidence. Following Oguz et al. (2022), both tables and infoboxes are linearized using simple templates. Specifically, we concatenate the properties and values from the table using the word "is". The entity name described by the infobox and the properties and values are strung together by a comma ",", forming a text string. Additionally, Wikipedia page titles are added at the beginning of the evidence for clearer information.
Types Avg. length Count Percentage
Text 19.86 5,916,596 57.74%
KG 11.40 2,214,854 21.61%
Table 20.32 1,043,105 10.18%
Infobox 11.05 1,072,440 10.47%
Sum 17.18 10,246,995 100.00%
To align with the standard IR task setup, we use automated scripts to label relevant evidence (golden labels) for each question. The relevance between the evidence and the question is of a boolean type (True/False). Specifically, if the entities in the evidence contain the answer to the question, the relevance is marked as True; otherwise, it is marked as False. Each question has at least one piece of evidence that provides the answer. The evidence retrieved for all questions in CompMix is combined into a heterogeneous knowledge pool, forming the corpus for the IR task. This corpus includes over 10 million pieces of evidence, covering knowledge about 137,808 different entities. Detailed statistics of CompMix-IR are presented in Table 1, and examples of linearized evidence and QA pair, as well as their annotation information examples, provided in Appendix A.
@@ 5.1 Baselines
Retrieval Scenario 1 (instruction $\bm{I}_{\text{All}}$) Retrieval Scenario 2 (instruction $\bm{I}_{\tau}$) Method Size Ins Hit@5 Hit@10 Hit@100 MRR@100 KG-Hit Text-Hit Table-Hit Info-Hit BM25 - 11.51 17.40 52.39 8.54 24.20 34.55 8.50 19.79 DPR 109M 24.89 36.32 78.76 17.51 49.13 63.68 15.63 41.57 Mpnet 109M 26.23 37.99 82.67 18.46 63.02 61.11 18.96 52.1 GTR-T5-base 110M 24.46 36.54 80.32 16.73 57.78 59.8 22.87 46.09 Contriever 109M 28.58 40.70 83.79 20.07 62.26 63.86 18.63 55.64 SimLM 109M 25.11 37.08 80.61 17.68 59.59 59.01 17.69 52.06 Instructor-base 110M 24.86 36.22 81.55 17.80 65.63 50.25 16.82 53.36 Instructor-large 336M 25.98 36.87 81.51 18.54 68.78 44.61 17.11 53.98 BGE 109M 26.66 39.04 84.15 19.40 68.42 57.96 22.58 56.58 BERT-finetuned 109M 24.46 35.38 78.51 17.04 57.63 54.67 17.55 48.41 UDT-retriever 109M 24.96 35.49 76.52 18.24 66.10 62.48 25.90 57.05 UniK-retriever 109M 30.68 43.42 85.20 21.22 67.40 63.21 26.74 56.04 UniHGKR-base 109M 32.38 45.55 85.75 22.57 75.43 70.30 41.24 66.21 Relative gain +5.54% +4.91% +0.65% +6.36% +9.67% +10.08% +54.23% +16.06%
Table 3: the experimental results for the two retrieval scenarios on CompMix-IR. The relative gain is calculated based on the performance of UniHGKR-base compared to the best baseline, highlighted by underlines.
Fine-tuned Baselines. We follow the verbalizer-retriever approach from UDT-QA Ma et al. (2022b) to fine-tune a BERT-base model, serving as the UDT retriever. Since UDT focuses on homogeneous textual representations of heterogeneous data, we replace $d_{i}$ with $t_{i}$ from the data-text pairs $\mathcal{D}$ during its training and evaluation, ensuring this model only interacts with the natural language corpus. This also means that in our experiments, we fine-tune the UDT-retriever baseline using exactly the same GPT-4o-mini synthesized data-text pairs $\mathcal{D}$ as utilized by our UniHGKR. For comparison, we also fine-tune a BERT-base model on the original CompMix-IR. Additionally, we finetune a DPR model using the UniK-QA method Oguz et al. (2022), serving as the UniK retriever. All fine-tuning uses the same positive and hard negative samples as UniHGKR. For baseline models lacking instruction-following capabilities, we input only the query across all retrieval scenarios to ensure optimal performance.
@@ 6.1 Main Results
Table 3 presents the retrieval performance of various models on the CompMix-IR test set. Our UniHGKR model outperforms all baselines in both scenarios, with a maximum relative improvement of 6.36% in scenario 1 and 54.23% in scenario 2, demonstrating its effectiveness in heterogeneous knowledge retrieval. Notably, powerful open-source retrievers like BGE (trained on over 200 million high-quality text pairs, Xiao et al. (2024)) only achieve an MRR@100 below 20.0 in scenario 1 and a Table-Hit of 22.58 in scenario 2, highlighting the challenges of our constructed benchmark. Although the UDT-retriever shows significant improvement over its counterpart model, BERT-finetuned, in Scenario 2, the improvement in Scenario 1 is minimal. Also, it is clearly inferior to our UniHGKR-base, which was trained on the same synthetic data from GPT-4o-mini. Moreover, UniK-retriever, fine-tuned the DPR model on CompMix-IR, performs well across several metrics but is suboptimal on structured data (like Table and Infobox) in scenario 2. In contrast, our UniHGKR shows the greatest improvements on metrics where existing methods struggle, particularly in retrieving structured knowledge in scenario 2. This indicates that our three-stage training approach not only creates an effective representation space for heterogeneous knowledge retrieval but also excels at following diverse user instructions.
@@ 6.2 Ablation Study
Retrieval Scenario 1 (instruction $\bm{I}_{\text{All}}$) Retrieval Scenario 2 (instruction $\bm{I}_{\tau}$) Method Hit@5 Hit@10 Hit@100 MRR@100 KG-Hit Text-Hit Table-Hit Info-Hit UniHGKR-base 32.38 45.55 85.75 22.57 75.43 70.30 41.24 66.21 w/o training stage 1 (pretrain) 29.78 42.80 84.88 21.54 72.97 68.60 34.01 60.06 w/o NL sentence during stage 1 31.30 44.36 85.46 21.83 74.07 69.01 40.13 65.16 w/o training stage 2 (alignment) 31.41 45.02 85.14 21.92 74.75 70.01 37.99 66.04 In the training stage 3 (finetune)   w/o type-preferred loss $\mathcal{L}_{\text{preferred}}$ 31.77 44.97 85.24 22.27 73.26 69.90 33.86 61.22   w/o instructions and $\mathcal{L}_{\text{preferred}}$ 31.98 44.39 85.24 22.18 68.20 65.59 29.20 58.65   w/o rewritten instructions 31.11 44.14 84.48 21.86 67.66 65.16 27.82 57.96
@@
Retrieval Scenario 1 (instruction $\bm{I}_{\text{All}}$) Retrieval Scenario 2 (instruction $\bm{I}_{\tau}$) Method Hit@5 Hit@10 Hit@100 MRR@100 KG-Hit Text-Hit Table-Hit Info-Hit UniHGKR-base 32.38 45.55 85.75 22.57 75.43 70.30 41.24 66.21 E5-mistral-7B 31.3 43.49 83.36 22.97 69.03 41.46 33.03 62.92 LLARA-passage 37.45 51.59 86.61 26.11 68.23 70.48 37.88 60.64 LLARA-finetuned 42.19 55.35 87.81 30.83 74.38 69.86 36.40 64.40 UniHGKR-7B 49.78 59.23 88.21 38.20 81.80 76.05 49.57 73.88 Relative gain +17.99% +7.01% +0.46% +23.91% +9.98% +7.90% +30.86% +14.72%
Table 5: Retrieval performances of UniHGKR-7B and LLM-based retrievers baselines. The relative gain is calculated based on the performance of UniHGKR-7B compared to the best baseline, highlighted by underlines.
In this subsection, we conduct ablation studies to examine the roles of different training stages and components in UniHGKR for heterogeneous knowledge retrieval. Table 4 presents the performance of various UniHGKR variants, obtained by removing specific components or a particular training stage. Results show that removing any training stage or component leads to a significant drop in performance. For retrieval scenario 1, training stage 1 (Unified Embedding Self-Supervised Pretraining) is crucial, while for scenario 2, both rewritten (paraphrased) instructions and instruction-aware type-preferred loss $\mathcal{L}_{\text{preferred}}$ are key. Removing them will result in a performance drop of up to 13.42 points in the Table-Hit metric. Additionally, we present some extra ablation studies in Appendix F, such as exploring the role of different instructions in retrieving from specific sources and the performance gains of different training stages in an unsupervised setting.